Which of this represent parallel lines


1)y-4x-4


2)y+Ax-4


3)y-4x-1


And Also Tell Procedure to solve this type of problem?

Grade:Upto college level

2 Answers

vikas askiitian expert
13 years ago

dear vikrant ,

 

 line 1 & 3 are parallel lines ...

 

y - 4x - 4 = 0                    L1

 

y -4x -1 = 0                       L2

comparing both of these eq by y-mx+c = 0 ...

 

m = slope = 4 for both these lines  so

 

if slopes are same then line are said to be parallel...

 

approve if u like this ans

Sudheesh Singanamalla
13 years ago

Dear vikrant ,

two lines a1x + b1y + c1 =0 and a2x + b2y + c2= 0 are said to be equal if

 

a1/a2 = b1/b2 ≠ c1/c2

where a,b,c are magnitudes (i.e numbers before x,y and c is constant)

so if we compare (1) and (3) we get it in terms of

a1/a2 = b1/b2 ≠ c1/c2

 

this is the simplest method of solving such type of problems, the long method is what Vikas Bhaiyya has done !

 

Please approve